# On discrete Sobolev inequalities for nonconforming finite elements under a semi-regular mesh condition

### v1

@@ Abstract
Mathematics Subject Classification (2020) 65N30 (primary); 65N15, 65N12, 46E35.
@@ 1 Introduction
This paper introduces discrete Sobolev inequalities applicable to nonconforming finite elements on semi-regular (anisotropic) meshes. Specifically, for the Crouzeix–Raviart (CR) finite element space and its discontinuous counterpart, we establish an $L^{q}-L^{p}$ ($q\leq p$) bound. Notably, the constants associated with this bound are influenced solely by the domain and the semi-regular parameter, remaining unaffected by the angles and aspect ratios of simplices. This extends the classical shape-regular theory (e.g. [2, 16]) to anisotropic partitions and provides a robust tool for stability and error analysis.
Let $\Omega\subset\mathbb{R}^{d}$, $d\in\{2,3\}$ be a bounded polyhedral domain. The Poisson problem is to find $u:\Omega\to\mathbb{R}$ such that
$\displaystyle-\varDelta u=f\quad\text{in $\Omega$},\quad u=0\quad\text{on $\partial\Omega$},$ (1.1)
where $f\in L^{2}(\Omega)$ is a given function. In [14, 7, 11], we considered the CR finite element type discretisation as follows. Find $u_{h}\in V_{h}$ such that
$\displaystyle a_{h}^{CR}(u_{h},v_{h}):=\int_{\Omega}\nabla_{h}u_{h}\cdot\nabla_{h}v_{h}dx+\sum_{F\in\mathcal{F}_{h}}\kappa_{F}\int_{F}\Pi_{F}^{0}[\![u_{h}]\!]\Pi_{F}^{0}[\![v_{h}]\!]ds$ $=\int_{\Omega}fv_{h}dx\quad\forall v_{h}\in V_{h},$ (1.2)
where $V_{h}$ is the CR finite element space defined in Section 2.3, $\kappa_{F}$ is a penalty parameter defined in [7, Section 2.4] and $\Pi_{F}^{0}$ is the $L^{2}$-projection defined in Section 2.4. We define an energy norm as
$\displaystyle|v_{h}|_{E}:=a_{h}^{CR}(v_{h},v_{h})^{\frac{1}{2}},\quad v_{h}\in V_{h}.$
Stability in the finite element method is the property that the norm of the discrete solution is controlled by a uniform constant, independent of the mesh and parameters, in the norm of the right-hand side (external force) data, which guarantees the reliability of the numerical solution. Substituting the finite element solution $u_{h}\in V_{h}$ for $v_{h}$ on the right-hand side of (1.2) yields
$\displaystyle|u_{h}|_{E}^{2}=\int_{\Omega}fv_{h}dx\leq\|f\|_{L^{2}(\Omega)}\|u_{h}\|_{L^{2}(\Omega)}.$
If the discrete Poincaré inequality:
$\displaystyle\|u_{h}\|_{L^{2}(\Omega)}\leq C^{P}_{dc}|u_{h}|_{E}$
holds, the following stability estimate is obtained.
$\displaystyle|u_{h}|_{E}\leq C^{P}_{dc}\|f\|_{L^{2}(\Omega)}.$
@@
Demonstrating the discrete Sobolev (Poincaré) inequality on anisotropic meshes presents significant challenges. Recent research has made notable progress in this area. The study in [7] established the discrete Poincaré inequality for the anisotropic weakly over-penalised symmetric interior penalty method. Following this, [9] extended these findings to a hybrid version of the method. Furthermore, [10] introduced the discrete Sobolev inequality on anisotropic meshes, specifically for the CR finite element method. The discrete Poincaré inequality as applied to Nitsche’s method on anisotropic meshes is detailed in [11]. These investigations utilise proofs based on the duality problem, necessitating the assumption of a convex domain $\Omega$ to substantiate the inequalities. This paper aims to give a new anisotropic discrete Sobolev inequality that removes the assumption that the domain is convex. Our results extend the classical shape-regular framework of Brenner (2003) to semi-regular meshes. For $q\mathop{\textgreater}p$, an $L^{p}-L^{2}$ estimate is available under a weak elliptic regularity (see Section 6 and [10, Lemma 4]).
@@
Section 2 establishes the notation and anisotropic preliminaries. Section 3 introduces semi-regular meshes and the Raviart–Thomas (RT) interpolation. Section 4 compiles the Bogovskiĭ operator, discrete integration-by-parts, and face estimates. Section 5 demonstrates the primary discrete $L^{q}-L^{p}$ inequality, while Section 6 provides concluding remarks.
@@
Throughout this paper, generic constants $c$ depend only on $d$, $\Omega$, $p$ and the semi-regular parameter; they are independent of $h$ (defined later), angles and aspect ratios of simplices. These values vary across different contexts. The notation $\mathbb{R}_{+}$ denotes a set of positive real numbers.
@@ 2 Preliminaries
Cross-references. See Section 2.1 for the jumps/averages and boundary treatment, Section 2.3 for the definitions of the CR/discontinuous-CR spaces and Section 2.4 for the discrete norms.
2.1 Mesh, Mesh faces, jumps and average
Let $\mathbb{T}_{h}=\{T\}$ be a simplicial mesh of $\overline{\Omega}$ composed of closed $d$-simplices:
$\displaystyle\overline{\Omega}=\bigcup_{T\in\mathbb{T}_{h}}T,$
where $h:=\max_{T\in\mathbb{T}_{h}}h_{T}$ and $h_{T}:=\mathop{\mathrm{diam}}(T)$. For simplicity, we assume that $\mathbb{T}_{h}$ is conformal; that is, $\mathbb{T}_{h}$ is a simplicial mesh of $\overline{\Omega}$ without hanging nodes. Throughout we assume $h\leq 1$.
Let $\mathcal{F}_{h}^{i}$ be the set of interior faces, and $\mathcal{F}_{h}^{\partial}$ be the set of faces on the boundary $\partial\Omega$. We set $\mathcal{F}_{h}:=\mathcal{F}_{h}^{i}\cup\mathcal{F}_{h}^{\partial}$. For any $F\in\mathcal{F}_{h}$, we define the unit normal $\bm{n}_{F}$ to $F$ as follows: (i) If $F\in\mathcal{F}_{h}^{i}$ with $F=T_{\natural}\cap T_{\sharp}$, $T_{\natural},T_{\sharp}\in\mathbb{T}_{h}$, $\natural>\sharp$, let $\bm{n}_{F}$ be the unit normal vector from $T_{\natural}$ to $T_{\sharp}$. (ii) If $F\in\mathcal{F}_{h}^{\partial}$, $\bm{n}_{F}$ is the unit outward normal $\bm{n}$ to $\partial\Omega$.
@@ 2.1 Mesh, Mesh faces, jumps and average
with a norm
@@
Especially, we write $H^{s}(\mathbb{T}_{h}):=W^{s,2}(\mathbb{T}_{h})$ with a norm
@@
Let $p\in[1,\infty)$. Let $\varphi\in W^{1,p}(\mathbb{T}_{h})$. Suppose that $F\in\mathcal{F}_{h}^{i}$ with $F=T_{+}\cap T_{-}$, $T_{+},T_{-}\in\mathbb{T}_{h}$, +>+>-. We set $\varphi_{+}:=\varphi{|_{T_{+}}}$ and $\varphi_{-}:=\varphi{|_{T_{-}}}$. We set two nonnegative real numbers $\omega_{T_{+},F}$ and $\omega_{T_{-},F}$ such that
@@
$\displaystyle[\![\varphi]\!]:=[\![\varphi]\!]_{F}:=\varphi_{+}-\varphi_{-},\quad\{\!\{\varphi\}\!\}_{\overline{\omega}}:=\{\!\{\varphi\}\!\}_{\overline{\omega},F}:=\omega_{T_{-},F}\varphi_{+}+\omega_{T_{+},F}\varphi_{-},\quad+>-.$
@@
$[\![{\bm{v}}\cdot{\bm{n}}]\!]:=[\![{\bm{v}}\cdot{\bm{n}}]\!]_{F}:={\bm{v}_{+}}\cdot{\bm{n}_{F}}-{\bm{v}_{-}}\cdot{\bm{n}_{F}},\quad+>-,$
$\{\!\{{\bm{v}}\}\!\}_{\omega}:=\{\!\{{\bm{v}}\}\!\}_{\omega,F}:=\omega_{T_{+},F}{\bm{v}_{+}}+\omega_{T_{-},F}{\bm{v}_{-}},\quad+>$-
@@ 2.2 Trace inequality
The trace inequality on anisotropic meshes discussed herein is of considerable importance to this study. The proof of this inequality is documented in several references. In this context, we adhere to the approach outlined by Ern and Guermond [4, Lemma 12.15]. It is noteworthy that, although Lemma 12.15 in [4] stipulates a shape-regular mesh condition, the condition is easily violated. For a simplex $T\subset\mathbb{R}^{d}$, let $\mathcal{F}_{T}$ be the collection of the faces of $T$. Let $|\cdot|_{d}$ denote the $d$-dimensional Hausdorff measure.
@@
Let $p\in[1,\infty]$. Let $T\subset\mathbb{R}^{d}$ be a simplex. There exists a positive constant $c$ such that for any $\bm{v}=(v^{(1)},\ldots,v^{(d)})^{\top}\in W^{1,p}(T)$, $F\in\mathcal{F}_{T}$, and $h$,
$\displaystyle\|\bm{v}\|_{L^{p}(F)^{d}}\leq c\ell_{T,F}^{-\frac{1}{p}}\left(\|\bm{v}\|_{L^{p}(T)^{d}}+h_{T}^{\frac{1}{p}}\|\bm{v}\|_{L^{p}(T)^{d}}^{1-\frac{1}{p}}|\bm{v}|_{W^{1,p}(T)^{d}}^{\frac{1}{p}}\right),$ (2.1)
where $\ell_{T,F}:=\frac{d!|T|_{d}}{|F|_{d-1}}$ denotes the distance of the vertex of $T$ opposite to $F$ to the face.
@@
Let $v\in W^{1,p}(T)$. Let $\bm{P}_{F}$ be the vertex of $T$ opposite to $F$. By the same argument of [4, Lemma 12.15], together with the fact that $|\bm{x}-\bm{P}_{F}|\leq h_{T}$ and $\frac{|F|_{d-1}}{|T|_{d}}=\frac{d!}{\ell_{T,F}}$, it holds that for $i\in\{1,\ldots,d\}$,
$\displaystyle\|v^{(i)}\|^{p}_{L^{p}(F)}$ $\leq\frac{d!}{\ell_{T,F}}\|v^{(i)}\|^{p}_{L^{p}(T)}+\frac{2(d-1)!}{\ell_{T,F}}h_{T}\|v^{(i)}\|^{p-1}_{L^{p}(T)}\|\bm{\nabla}v^{(i)}\|_{L^{p}(T)^{d}}.$
Using the Cauchy–Schwarz inequality, we obtain the target inequality together with Jensen’s inequality. ∎

### v2

@@ Abstract
We derive a discrete $L^{q}-L^{p}$ Sobolev inequality tailored for the Crouzeix–Raviart and discontinuous Crouzeix–Raviart finite element spaces on anisotropic meshes in both two and three dimensions. Subject to a semi-regular mesh condition, this discrete Sobolev inequality is applicable to all pairs $(q,p)$ that align with the local Sobolev embedding, including scenarios where $q\leq p$. Importantly, the constant is influenced solely by the domain and the semi-regular parameter, ensuring robustness against variations in aspect ratios and interior angles of the mesh. The proof employs an anisotropy-sensitive trace inequality that leverages the element height, a two-step affine/Piola mapping approach, the stability of the Raviart–Thomas interpolation, and a discrete integration-by-parts identity augmented with weighted jump/trace terms on faces. This Sobolev inequality serves as a mesh-robust foundation for the stability and error analysis of nonconforming and discontinuous Galerkin methods on highly anisotropic meshes.
Keywords Discrete Sobolev inequalies ;Anisotropic (semi-regular) meshes; Crouzeix–Raviart finite elements; Nonconforming finite elements.
Mathematics Subject Classification (2020) 65N30 (primary); 65N15, 65N12, 46E35
@@ 1 Introduction
This paper introduces discrete Sobolev inequalities applicable to nonconforming finite elements on semi-regular (anisotropic) meshes. Specifically, for the Crouzeix–Raviart (CR) finite element space and its discontinuous counterpart, we establish an $L^{q}-L^{p}$ ($q\leq p$) bound. Significantly, the constants associated with this bound are exclusively determined by the dimension, the domain, the exponents $p$ and $q$, and the semi-regular mesh parameter. These constants remain unaffected by the mesh size, as well as the aspect ratios and interior angles of the simplices. This extends the classical shape-regular theory (e.g. [24, 25, 23, 6, 22]) to anisotropic partitions and provides a robust tool for stability and error analysis.
Let $\Omega\subset\mathbb{R}^{d}$, $d\in\{2,3\}$, be a bounded Lipschitz polyhedral domain. The limitation to $d\in\{2,3\}$ arises from the application of anisotropic Raviart–Thomas (RT) interpolation error estimates on semi-regular meshes as detailed in [12]. These estimates are derived through a two-step affine-transformation technique (combined with the associated Piola transform for RT fields) that employs an explicit curvilinear parametrisation of the geometry of triangles and tetrahedra. Extending this framework to $d\geq 4$ would necessitate a corresponding higher-dimensional geometric parametrisation (or an alternative coordinate-free reformulation) along with new anisotropic bounds for the RT facet moments; such results are currently not available. We leave this extension as a challenging avenue for future research. In contrast, the Bogovskiĭ-type right inverse of the divergence, as well as the Poincaré and trace inequalities used in this paper, are applicable in any spatial dimension. As a representative model problem that motivates the discrete Sobolev inequalities proved below, we recall the Poisson problem: find $u:\Omega\to\mathbb{R}$ such that
$\Delta u=f\quad\text{in }\Omega,\qquad u=0\quad\text{on }\partial\Omega,$ (1.1)
where $f\in L^{2}(\Omega)$ is a given function.
For nonconforming and discontinuous Galerkin discretisations of (1.1), discrete Poincaré–Sobolev inequalities of the type studied in this paper lead to mesh-robust stability bounds of the form
$\displaystyle|u_{h}|_{E}\leq C^{\mathrm{stab}}\|f\|_{L^{2}(\Omega)}.$
where $|\cdot|_{E}$ denotes the natural energy norm on the discrete space and $C^{\mathrm{stab}}$ is independent of the mesh size $h$ and of the aspect ratios and interior angles of the simplices; see, for example, the CR-type schemes analysed in [13, 17]. This provides one of the main motivations for deriving discrete $L^{q}-L^{p}$ Sobolev inequalities under a semi-regular mesh condition in the present work. Discrete Sobolev embeddings of the $L^{q}-L^{p}$ type are a fundamental tool in the analysis of nonlinear problems. For example, they are integral to the stability analysis of nonconforming discretisations of $p$-Laplacian-type models, and more broadly, problems involving nonlinear lower-order terms. In such contexts, it is essential to have a mesh-robust embedding estimate $\|u_{h}\|_{L^{q}(\Omega)}\leq C^{\mathrm{pLap}}|u_{h}|_{W^{1,p}(\mathbb{T}_{h})}$, where $C^{\mathrm{pLap}}$ remains independent of the mesh size $h$, as well as the aspect ratios and interior angles of the simplices, to control the nonlinear terms. Similarly, in the case of the stationary incompressible Navier–Stokes equations, these embeddings are employed to estimate the convection term, specifically the associated trilinear form, in terms of broken $W^{1,p}$ seminorms.
@@
Discrete Poincaré- and Sobolev-type inequalities for nonconforming CR-type finite element spaces have, in fact, a longer history. Early instances of such estimates can already be found in the works of Temam and Thomas in the context of stationary Navier–Stokes equations and mixed or nonconforming finite element approximations on shape-regular meshes; see, for example, [24, 25, 23]. [6] later provides a systematic treatment of Poincaré–Friedrichs inequalities for piecewise $H^{1}$ functions on shape-regular partitions including nonconforming partitions with hanging nodes and edges.
Anisotropic finite elements, under mesh assumptions related to the classical maximum-angle condition, have been extensively investigated. In [1], Apel established a comprehensive framework for anisotropic finite elements, offering precise local interpolation estimates and geometric tools for addressing anisotropic simplicial meshes. Building upon this framework, [2] conducted an in-depth analysis of CR-type finite elements on anisotropic meshes in three dimensions. They derived optimal-order finite element error estimates for a nonconforming approximation of the Poisson problem by integrating anisotropic interpolation bounds with consistency error estimates. To the best of our knowledge, neither [1] nor [2] explicitly formulates global discrete $L^{q}-L^{p}$ Sobolev or Poincaré inequalities for CR-type spaces as a primary focus.
Demonstrating the discrete Sobolev (Poincaré) inequality on anisotropic meshes presents significant challenges. Recent research has made notable progress in this area. For the discontinuous Galerkin-type finite element space underlying an anisotropic weakly over-penalised symmetric interior penalty (WOPSIP) formulation of the Stokes equations, [13] established a discrete Poincaré inequality on anisotropic semi-regular meshes. For the CR finite element space, [16] derived discrete $L^{q}-L^{p}$ Sobolev inequalities on anisotropic meshes under a semi-regular mesh condition and the weak elliptic regularity assumption. In particular, when $p=q=2$, the proof in [16] uses a duality argument posed on a convex domain; see also the references therein for related variants such as hybrid interior penalty and Nitsche-type formulations. These investigations rely on duality-based proofs, for which the convexity of $\Omega$ is crucial in order to validate the inequalities. The present paper develops an alternative approach based on the Bogovskiǐ operator and RT interpolation, which applies to general bounded Lipschitz polyhedral domains and therefore removes the convexity restriction. In the context of the Hilbert space where $p=q=2$, related Poincaré–Friedrichs-type estimates concerning general decompositions under appropriate trace inequalities can be traced back at least to [3, Lemma 2.1]. Our findings also address the scenario where $p=q=2$, and we remove the convexity restrictions typically associated with duality and elliptic regularity arguments by utilising the Bogovskiǐ/RT-based approach on bounded Lipschitz polyhedral domains.
Our findings expand the application of discrete Poincaré–Friedrichs and Sobolev inequalities to piecewise $H^{1}$ functions on shape-regular partitions. This expansion includes nonconforming partitions with hanging nodes and edges, as described in [6], and further extends these inequalities to anisotropic semi-regular simplicial meshes. The semi-regular mesh condition employed in this paper is expressed in terms of a geometric flatness parameter; for tetrahedral meshes, it was shown in [18] that this semi-regular condition is in fact equivalent to the classical maximum-angle condition. For $q\mathop{\textgreater}p$, an $L^{p}-L^{2}$ estimate is available under a weak elliptic regularity (see Section 6 and [16, Lemma 4]).
@@
Section 2 establishes the notation and anisotropic preliminaries. Section 3 introduces semi-regular meshes and the RT interpolation. Section 4 compiles the Bogovskiĭ operator, discrete integration-by-parts, and face estimates. Section 5 demonstrates the primary discrete $L^{q}-L^{p}$ inequality, while Section 6 provides concluding remarks.
@@
Throughout this paper, we denote by $c$ generic constants, independent of $h$ (defined later) and the angles and aspect ratios of simplices; unless specified otherwise, all constants $c$ are bounded if the maximum angle is bounded. These values vary across different contexts. The notation $\mathbb{R}_{+}$ denotes a set of positive real numbers.
@@ 2 Preliminaries
In this section, we collect the notation, mesh assumptions and anisotropic preliminaries used in the subsequent analysis.
2.1 Mesh, mesh faces, jumps and averages
Let $\mathbb{T}_{h}$ be a simplicial mesh of $\overline{\Omega}$ composed of closed $d$-simplices. Let $N_{e}:=\#\mathbb{T}_{h}$ denote the number of elements of the mesh and write
$\displaystyle\mathbb{T}_{h}=\{T_{j}\}_{j=1}^{N_{e}},\quad\overline{\Omega}=\bigcup_{j=1}^{N_{e}}T_{j}.$
For each $j\in\{1,\ldots,N_{e}\}$, we set
$\displaystyle h_{T_{j}}:=\mathop{\mathrm{diam}}(T_{j}),\quad h:=\max_{1\leq j\leq N_{e}}h_{T_{j}}.$
For simplicity, we assume that $\mathbb{T}_{h}$ is conforming; that is, $\mathbb{T}_{h}$ is a simplicial mesh of $\overline{\Omega}$ with no hanging nodes. Throughout this paper, we assume $0<h\leq 1$.
Remark 2.1 (Conforming meshes).
The assumption of a conforming mesh (i.e., the absence of hanging nodes) is made solely to simplify the face-based notation and constructions. In particular, it ensures that each interior face $F$ is a single (d1)(d-1)-simplex shared by exactly two elements, so that the unit normal and the jump and average operators on $F$ are unambiguous. This assumption also streamlines the definition of global face-based finite element spaces, such as the CR-type nonconforming spaces and the $H(\mathop{\mathrm{div}})$-conforming RT space, where face degrees of freedom (including normal-flux moments for RT) are matched consistently across each interior face.
Meshes with hanging nodes can, in principle, be treated by working on the induced subdivision of the mesh skeleton into subfaces and by redefining the face-based operators and matching conditions accordingly; for CR-type settings, such constructions are classical, see e.g., [7, Section 2.7]. For the $H(\mathop{\mathrm{div}})$-conforming RT space, however, hanging faces typically require additional design choices and constraints to enforce consistent matching of normal-flux moments across a coarse face and its subfaces (while preserving the standard commuting and approximation properties). To avoid this additional technical overhead, we restrict ourselves to conforming meshes in the present paper.
@@ 2.1 Mesh, mesh faces, jumps and averages
equipped with the broken Sobolev seminorm
@@
In particular, we write $H^{s}(\mathbb{T}_{h}):=W^{s,2}(\mathbb{T}_{h})$ and
@@
We remark that in the present paper, we only use the broken Sobolev space $W^{s,p}(\mathbb{T}_{h})$ with the integer order $s=1$, i.e. $W^{1,p}(\mathbb{T}_{h})$ and $H^{1}(\mathbb{T}_{h})=W^{1,2}(\mathbb{T}_{h})$. In particular, the Bogovskiĭ-type right-inverse of the divergence in Lemma 4.1 is applied only in the classical setting $L^{r}_{0}(\Omega)\to W^{1,r}_{0}(\Omega)^{d}$ for $r\in(1,\infty)$. Thus, our use of the Bogovskiĭ operator is restricted to the standard integer-order setting, and no results for fractional-order Sobolev spaces are needed.
Furthermore, throughout the present paper, we restrict the exponent to $p\in[1,\infty)$. This is the natural range in which we use the trace theorem for $W^{1,p}$ functions on Lipschitz domains (see, e.g., [9, Theorem 3.10]), in order to give a precise meaning to face traces and thus to jumps and averages. Since our discrete $L^{q}-L^{p}$ Sobolev inequalities in Section 5 are also formulated only for finite exponents $p$ and $q$, the endpoint case $p=\infty$ is not needed in the sequel.
Let $\mathcal{F}_{h}^{i}$ be the set of interior faces, and $\mathcal{F}_{h}^{\partial}$ be the set of faces on the boundary $\partial\Omega$. We set $\mathcal{F}_{h}:=\mathcal{F}_{h}^{i}\cup\mathcal{F}_{h}^{\partial}$. For any $F\in\mathcal{F}_{h}$, we define the unit normal $\bm{n}_{F}$ to $F$ as follows: (i) If $F\in\mathcal{F}_{h}^{i}$ with $F=\partial T_{+}\cap\partial T_{-}$, $T_{+},T_{-}\in\mathbb{T}_{h}$, +>+>-, let $\bm{n}_{F}$ be the unit normal vector from $T_{+}$ to $T_{-}$. (ii) If $F\in\mathcal{F}_{h}^{\partial}$, $\bm{n}_{F}$ is the unit outward normal $\bm{n}$ to $\partial\Omega$. We fix once and for all a local ordering of the neighbouring elements and use the symbol ”+>+>-” to indicate that $T_{+}$ is the first element and $T_{-}$ the second one in this ordering.
Let $p\in[1,\infty)$. Let $\varphi\in W^{1,p}(\mathbb{T}_{h})$. Suppose that $F\in\mathcal{F}_{h}^{i}$ with $F=\partial T_{+}\cap\partial T_{-}$, $T_{+},T_{-}\in\mathbb{T}_{h}$. For such a face, we set $\varphi_{+}:=\varphi{|_{T_{+}}}$ and $\varphi_{-}:=\varphi{|_{T_{-}}}$. We choose two nonnegative real numbers $\omega_{T_{+},F}$ and $\omega_{T_{-},F}$ with
@@
$\displaystyle[\![\varphi]\!]:=[\![\varphi]\!]_{F}:=\varphi_{+}-\varphi_{-},\quad\{\!\{\varphi\}\!\}_{\overline{\omega}}:=\{\!\{\varphi\}\!\}_{\overline{\omega},F}:=\omega_{T_{-},F}\varphi_{+}+\omega_{T_{+},F}\varphi_{-}.$
@@
$[\![{\bm{v}}\cdot{\bm{n}}]\!]:=[\![{\bm{v}}\cdot{\bm{n}}]\!]_{F}:={\bm{v}_{+}}\cdot{\bm{n}_{F}}-{\bm{v}_{-}}\cdot{\bm{n}_{F}},$
$\{\!\{{\bm{v}}\}\!\}_{\omega}:=\{\!\{{\bm{v}}\}\!\}_{\omega,F}:=\omega_{T_{+},F}{\bm{v}_{+}}+\omega_{T_{-},F}{\bm{v}_{-}}$
@@
This clarifies the meaning of ”+>+>-” wherever it is used in the manuscript.
@@ 2.2 Trace inequality
The trace inequality on anisotropic meshes discussed herein is of considerable importance to this study. The proof of this inequality is documented in several references. In this context, we adhere to the approach outlined by Ern and Guermond [9, Lemma 12.15]. It is noteworthy that, although Lemma 12.15 in [9] stipulates a shape-regular mesh condition, the condition is easily violated. For a simplex $T\subset\mathbb{R}^{d}$, let $\mathcal{F}_{T}$ be the collection of the faces of $T$. Let $|\cdot|_{d}$ denote the $d$-dimensional Hausdorff measure.
@@
Let $p\in[1,\infty]$. Let $T\subset\mathbb{R}^{d}$ be a simplex. There exists a positive constant $C^{Tr}(d,p)$} such that for any $\bm{v}=(v^{(1)},\ldots,v^{(d)})^{\top}\in W^{1,p}(T)$, $F\in\mathcal{F}_{T}$, and $h$,
$\displaystyle\|\bm{v}\|_{L^{p}(F)^{d}}\leq{C^{Tr}(d,p)}\ell_{T,F}^{-\frac{1}{p}}\left(\|\bm{v}\|_{L^{p}(T)^{d}}+h_{T}^{\frac{1}{p}}\|\bm{v}\|_{L^{p}(T)^{d}}^{1-\frac{1}{p}}|\bm{v}|_{W^{1,p}(T)^{d}}^{\frac{1}{p}}\right),$ (2.1)
where $\ell_{T,F}:=\frac{d|T|_{d}}{|F|_{d-1}}$ denotes the distance of the vertex of $T$ opposite to $F$ to the face. Here, the constant $C^{Tr}(d,p)$ depends only on the dimension $d$ and the exponent $p$ and is independent of $T$, $F$, the mesh size $h_{T}$ and the semi-regular parameter $\gamma_{0}$.
@@
The estimate is a direct consequence of the anisotropic trace inequality on simplices proved in [9, Lemma 12.15]. More precisely, applying [9, Lemma 12.15] to each component $v^{(i)}$ of $\bm{v}=(v^{(1)},\ldots,v^{(d)})^{\top}$ and expressing the resulting bound in terms of the height $\ell_{T,F}=\frac{d|T|_{d}}{|F|_{d-1}}$ yields the scalar version of (2.1) for $v^{(i)}$. Summing over $i=1,\ldots,d$ and using the Cauchy–Schwarz inequality then gives (2.1) with a constant depending only on $d$ and $p$. We refer to [9, Lemma 12.15] for the detailed scalar proof.